Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Carl Edelmann GmbH&Co. KG Roland Combosch Sales Manager Shooting Sport Targets P.O. Box 12 45 D - 89502 Heidenheim Germany

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Catalog Request  21st June 2001

A Chara,

I am writing as the secretary of the Dublin University Rifle Club. We are a college club numbering approximately three to four hundred members per year. We use an appreciatable amount of targets per year, for 10 metre air rifle and 50 metre prone rifle, both ISSF standard, and NSRA standard 25 yard prone rifle.

I would like to request a current catalog, a current price list, and any details about ordering directly from Edelmann to Dublin, Ireland.

Misé le meas,

Mark Dennehy, DURC Secretary